OpenStudy (anonymous):

an object lying on the ground is moved with constant acceleration 2 m/s^2. what is the velocity 5 seconds later?

OpenStudy (abhisar):

Here, u=0 a=2 t=5 v=? v=u+at => v=0+2*5 = 10m/s